the supplement of an angle is four times the complement of the angle find the measure of half the supplement of the angle

Answer by math_tutor2020(3817)


x = the unknown angle
180-x = the supplement of this angle
90-x = the complement of the original angle
Supplementary angles add to 180 degrees.
Complementary angles add to 90 degrees.

The instructions mention that the supplement is 4 times the complement,
supplement = 4*complement
180-x = 4*(90-x)
180-x = 360-4x
-x+4x = 360-180
3x = 180
x = 180/3
x = 60 is the original angle
180-x = 180-60 = 120 is the supplement, which cuts in half to 60 degrees, so that is the final answer.